# Residual Connections Harm Abstract Feature Learning in Masked Autoencoders

### v2

@@ Abstract
We demonstrate that adding a weighting factor to decay the strength of identity shortcuts within residual networks substantially improves semantic feature learning in the state-of-the-art self-supervised masked autoencoding (MAE) paradigm. Our modification to the identity shortcuts within a VIT-B/16 backbone of an MAE boosts linear probing accuracy on ImageNet from 67.8% to 72.7%. This significant gap suggests that, while residual connection structure serves an essential role in facilitating gradient propagation, it may have a harmful side effect of reducing capacity for abstract learning by virtue of injecting an echo of shallower representations into deeper layers. We ameliorate this downside via a fixed formula for monotonically decreasing the contribution of identity connections as layer depth increases. Our design promotes the gradual development of feature abstractions, without impacting network trainability. Analyzing the representations learned by our modified residual networks, we find correlation between low effective feature rank and downstream task performance.
@@ 1 Introduction
Figure 1: We propose decayed identity shortcuts, a variant of standard residual connections, to facilitate self-supervised feature representation learning. Our design introduces a depth-dependent scaling factor to shortcuts in the residual network, which modulates the contribution of preceding layers and fosters greater abstraction in deeper layers. Implementing our design in a Masked Autoencoder (MAE) [1], we show a substantial improvement in feature quality when benchmarking linear probing accuracy on the ImageNet-100 dataset, achieved without the introduction of additional trainable parameters. We further validate our design in terms of feature abstraction by demonstrating that it reduces the effective rank (defined in Section 5) of the output features.
@@
Residual networks (ResNets) [2] define a connection structure that has achieved near-universal adoption into modern architectures for deep learning. At the time of their development, supervised learning (e.g., ImageNet [3] classification) was the driving force behind the evolution of convolutional neural network (CNN) architectures. Residual networks solved a key issue: CNNs constructed of more than approximately 20 convolutional layers in sequence became difficult to train, leading to shallower networks outperforming deeper ones, unless additional techniques, such as auxiliary outputs [4] or batch normalization [5], were employed. Both ResNets, and their predecessor, highway networks [6] provide elegant solutions to this trainability problem by endowing the network architecture with alternative shortcut pathways along which to propagate gradients. Highway networks present a more general formulation that modulates these shortcut connections with learned gating functions. However, given their sufficient empirical effectiveness, the simplicity of ResNet’s identity shortcuts (residual connections) makes them a preferred technique.
@@
While solving the gradient propagation issue, residual connections impose a specific functional form on the network; between residual connections, each layer (or block of layers) learns to produce an update slated to be added to its own input. This incremental functional form may influence the computational procedures learned by the network [7]. Alternatives to residual and highway networks exist that do not share this functional form, but implement other kinds of skip-connection scaffolding in order to assist gradient propagation [8, 9, 10]. Thus, shortcut pathways, rather than a specific form of skip connection, are the essential ingredient to enable the training of very deep networks. Nevertheless, nearly all modern large-scale models, including those based on the transformer architecture [11] incorporate the standard identity shortcut residual connection.
@@
This design choice holds, even as deep learning has shifted into an era driven by self-supervised training. The shift to self-supervision brings to the forefront new learning paradigms, including those based on contrastive [12, 13, 14, 15, 16], generative [17, 18, 19, 20, 21, 22], and autoencoding [23, 1, 24] objectives. Many systems in the generative and autoencoding paradigms rely on “encoder-decoder” architectures, often styled after the original U-Net [25], which contains additional long-range shortcuts between corresponding layers in mirrored symmetry about a central bottleneck. With representation learning as a goal, one typically desires that the middle bottleneck layer produce a feature embedding reflecting abstract semantic properties. The interaction of skip-connection scaffolding for gradient propagation with encoder-decoder architectures, self-supervised training objectives, and bottleneck representations has not been carefully reconsidered. This is a worrisome oversight, especially since even in the supervised setting with standard classification architectures, prior work suggests that unweighted identity shortcuts may be a suboptimal design decision [26, 27].
@@
Intuitively, identity shortcut connections may not be entirely appropriate for capturing high-level, semantic representations as they directly inject low-level, high-frequency details of inputs into outputs, potentially compromising feature abstraction. Our work explores this issue within the Masked Autoencoder (MAE) framework of He et al. [1], one of the leading paradigms for self-supervised image representation learning. We demonstrate that identity shortcuts significantly harm semantic feature learning in comparison to an alternative we propose: gradually decay the weight of the identity shortcut over the depth of the network, thereby reducing the information flowing through it (Figure 1). With increasing layer depth, our approach facilitates a smooth transition from a residual to a feed-forward architecture, while maintaining sufficient connectivity to train the network effectively. Unlike prior work on learned gating [6] or reweighting [26] mechanisms for residual connection, our method is a forced decay scheme governed by a single hyperparameter; Section 3 provides details.
@@
A parallel motivation for our design stems from Huh et al. [28], who show that features from residual blocks have higher rank than those produced by comparative feed-forward blocks. The smooth transition between residual and feed-forward behavior induced by our decay scheme should regularize deeper features toward exhibiting low-rank characteristics. Section 5 experimentally explores the correlation between our decayed identity shortcuts and low-rank feature representations. Our contributions are as following:
@@
    We introduce decayed identity shortcuts, a novel simplified mechanism governed by a single hyper-parameter, which enhances semantic feature abstraction in masked autoencoders.
@@
    Our novel design within an MAE yields a substantial performance boost in linear probing on ImageNet-1K [3] (72.7% from a baseline 67.8%).
@@
    Ablation studies on ImageNet-100 show that smaller models equipped with our decayed identity shortcuts can outperform larger ones equipped with standard residual connections. In particular, a VIT-S/16 model [29] with our design outperforms a baseline VIT-B/16 (78.5% vs. 76.5%).
@@ 2 Related Work
Self-supervised representation learning. Recent advancements [30, 31, 22, 32, 33, 34] in deep learning follow a common scaling law, in which a model’s performance consistently improves with its capacity and the size of the training data. This effect can be observed in large language models (LLMs), which are trained on vast amounts of internet text, enabling them to perform some tasks at human level [35] and exhibit remarkable zero-shot capabilities [36]. These models are trained using next-token-prediction, allowing them to be trained without labeled data. In contrast, the progress of this scaling law in computer vision has largely depended on annotated data. For instance, the Segment Anything model [31] leverages 1 billion human-annotated masks, and state-of-the-art image generators [34] require training on huge datasets of text-image pairs [37]. However, the vast volume of unlabeled visual data and desire for continued scaling motivates a transition to self-supervised learning.
@@
At present, two families of approaches to self-supervised visual representation learning appear particularly promising: contrastive learning [12, 13, 14, 15, 16], which trains a discriminative model to maximize mutual information across image augmentations, and masked image modeling [38, 1, 39], which trains a generative model to predict occluded pixels. Hybrid approaches [40, 41, 24] combine both methods. Despite advancements, neither approach has demonstrated the same scalability [42] as seen in LLMs. This challenge is additional motivation for reconsidering the foundations of self-supervised network architectures.
@@
Residual and skip-connection architectures. Highway networks [7] first propose an additive skip connection structure to provide a scaffolding for gradient propagation when training very deep (e.g., 100 layer) networks. Motivated by the gating mechanisms within LSTMs [43], this solution uses learned gating functions to weight each combination of identity and layer output branches. Residual networks [2] are a simplification that removes these learned coefficients. DenseNet [9] and FractalNet [8] demonstrate that access to gradient paths of multiple lengths are the core requirement of training scaffolding, by introducing skip-connection structures with other functional forms. DenseNet utilizes feature concatenation instead of addition, while FractalNet imposes a recursive tree-like architecture combining subnetworks of multiple depths.
@@
Zhu et al. [44] explore variants of ResNets and DenseNets with fewer points of combination between different internal paths, demonstrating that a sparser scaffolding structure may be more robust as network depth increases to thousands of layers. Savarese and Figueiredo [26] add a scalar gating functional to the layer output in residual networks, yielding a hybrid design between residual and highway networks; learning this scalar gating provides a consistent benefit to classification accuracy. Fischer et al. [27] develop a weighting scheme for residual connections based upon a sensitivity analysis of signal propagation within a ResNet. To date, none of these potential improvement have seen broad adoption.
@@
Low rank bias in neural networks. Over-parameterized neural networks exhibit surprising generalization capabilities, a finding seemingly in contradiction with classical machine learning theory [45]. This phenomenon implies the existence of some form of implicit regularization that prevents the model from overfitting. From the perspective of neural network parameterizations, Arora et al. [46] suggest that linear models with more layers tend to converge to minimal norm solutions. In the context of CNNs, Huh et al. [28] demonstrates that stacking more feed-forward layers compels the model to seek solutions of a lower rank, and Jing et al. [47] reinforce this finding by adding more layers to an Autoencoder’s bottleneck, thereby creating a representation bottleneck. In vision transformers, Geshkovski et al. [48] examine the connection between attention blocks and mean-shift clustering [49], showing that repeated attention operations result in low-rank outputs. Moreover, Dong et al. [50] reveal that eliminating the shortcut connection from residual attention blocks causes features to degenerate to rank 1 structures doubly exponentially. From a different perspective, recent works [51, 52, 53] show training algorithms implicitly induce low-rank behavior in neural networks. Radhakrishnan et al. [53] study the dimensionality reduction behavior of a recursive feature machine [51] and effectively verify performance on low-rank matrix recovery.
@@ 3 Method
Prior works show that deeper feed-forward architectures have an inductive bias towards producing low-rank feature maps, while ResNets do not display the same behavior [28]. However, despite this bias, deeper feed-forward architectures are typically less effective and generalize worse than ResNets [2]. We aim to combine the properties of both feed-forward networks and ResNets, using the low-rank prior to enhance the abstraction capability of the network while maintaining the core benefits of the residual block, including stable training and the capacity to construct deeper models.
@@ 3.1 Decayed Identity Shortcuts
where ${\bm{x}}_{l}\in\mathbb{R}^{d}$ represents the output from the preceding layer, and $f_{{\bm{\theta}}_{l}}$ denotes the transformation applied at the current layer. Although it is widely known that pure feed-forward architectures are susceptible to vanishing gradients when building deeper models, Huh et al. [28] demonstrates that feed-forward modules offer implicit structural regularization, enabling deep models to generate abstract representations at bottlenecks.
Residual connections. To address the optimization problem of vanishing gradients in deeper neural networks, ResNets [2] construct each layer as a residual function, resulting in a modification to Eq. 1:
@@
This design builds shortcuts from input to output, allowing gradient magnitude to be preserved regardless of the depth of the model. However, a consequence of this design is that the output stays close to the input in practice [7], defeating the need to construct complex transformations over depth. The same phenomenon is also observed in highway networks [6], which adopt learnable gates $H_{\phi}({\bm{x}})\in[0,1]^{d}$ in both the residual and skip branches: ${\bm{x}}_{l+1}=H_{\phi}({\bm{x}}_{l})\cdot{\bm{x}}_{l}+(1-H_{\phi}({\bm{x}}_{l}))\cdot f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})$. Although this flexible design allows the model to build the abstraction level over depths, similar to feedforward networks, Srivastava et al. [54] finds $H_{\phi}\approx 1$ for most units, suggesting the model prefers copying the input.
Decayed identity shortcuts for unsupervised representation learning. Setting aside the optimization benefits brought by residual connections, we rethink the role of the residual connections from the viewpoint of representation learning. Abstraction can be viewed as invariance to local changes of input and is crucial to the disentanglement of the feature space [55]. Prior work suggests that a shortcut path of residual connections tends to preserve high-frequency fine-grained input information [7], resulting in decreased feature abstraction. We hypothesize that this lack of abstraction harms the capability of the model to learn meaningful low-level features and that ensuring an abstract structure in the deeper layers of the neural network will help improve representation learning, especially for unsupervised tasks that often use indirect proxy objectives, such as pixel-wise reconstruction loss. Motivated by this hypothesis, we propose to downweight the contribution from the shortcut path:

### v3

# Formatting Instructions for ICLR 2025 Conference Submissions

@@ Abstract
We show that introducing a weighting factor to reduce the influence of identity shortcuts in residual networks significantly enhances semantic feature learning in generative representation learning frameworks, such as masked autoencoders (MAEs) and diffusion models. Our modification improves linear probing accuracy for both, notably increasing ImageNet accuracy from 67.8% to 72.7% for MAEs with a VIT-B/16 backbone, while also boosting generation quality for diffusion models. This significant gap suggests that, while residual connection structure serves an essential role in facilitating gradient propagation, it may have a harmful side effect of reducing capacity for abstract learning by virtue of injecting an echo of shallower representations into deeper layers. We ameliorate this downside via a fixed formula for monotonically decreasing the contribution of identity connections as layer depth increases. Our design promotes the gradual development of feature abstractions, without impacting network trainability. Analyzing the representations learned by our modified residual networks, we find correlation between low effective feature rank and downstream task performance.
Figure 1: We design decayed identity shortcuts (Figure 2), a variant of residual connections, to facilitate self-supervised representation learning. Compared to standard residual connections, our approach yields superior abstract semantic features (left, visualized using Zhang et al. (2024)’s approach), whose leading components pop out object instances and classes. Quantitative evaluation shows our architecture encourages lower feature rank and achieves a substantial increase in linear probing accuracy for both MAE and diffusion models, along with enhanced generation quality for diffusion models (right). These improvements require no additional learnable parameters.
@@ 1 Introduction
Residual networks (ResNets) (He et al., 2016) define a connection structure that has achieved near-universal adoption into modern architectures for deep learning. At the time of their development, supervised learning (e.g., ImageNet (Deng et al., 2009) classification) was the driving force behind the evolution of convolutional neural network (CNN) architectures. Residual networks solved a key issue: CNNs constructed of more than approximately 20 convolutional layers in sequence became difficult to train, leading to shallower networks outperforming deeper ones, unless additional techniques, such as auxiliary outputs (Szegedy et al., 2015) or batch normalization (Ioffe & Szegedy, 2015), were employed. Both ResNets, and their predecessor, highway networks (Srivastava et al., 2015a) provide elegant solutions to this trainability problem by endowing the network architecture with alternative shortcut pathways along which to propagate gradients. Highway networks present a more general formulation that modulates these shortcut connections with learned gating functions. However, given their sufficient empirical effectiveness, the simplicity of ResNet’s identity shortcuts (residual connections) makes them a preferred technique.
Figure 2: Our decayed identity shortcuts introduce a depth-dependent scaling factor to shortcuts in a residual network, thereby modulating the contribution of preceding layers and fostering greater abstraction in deeper layers. A simple schema for controlling decay factor $\alpha$ suffices to improve feature learning in both MAEs and diffusion models, as well as diffusion model generation quality.
@@
While solving the gradient propagation issue, residual connections impose a specific functional form on the network; between residual connections, each layer (or block of layers) learns to produce an update slated to be added to its own input. This incremental functional form may influence the computational procedures learned by the network (Greff et al., 2017). Alternatives to residual and highway networks exist that do not share this functional form, but implement other kinds of skip-connection scaffolding in order to assist gradient propagation (Larsson et al., 2017; Huang et al., 2017; Zhu et al., 2018). Thus, shortcut pathways, rather than a specific form of skip connection, are the essential ingredient to enable the training of very deep networks. Nevertheless, nearly all modern large-scale models, including those based on the transformer architecture (Vaswani et al., 2017) incorporate the standard identity shortcut residual connection.
@@
This design choice holds, even as deep learning has shifted into an era driven by self-supervised training. The shift to self-supervision brings to the forefront new learning paradigms, including those based on contrastive (Wu et al., 2018; He et al., 2020; Chen et al., 2020; Caron et al., 2021; Grill et al., 2020), generative (Goodfellow et al., 2014; Karras et al., 2021; Ho et al., 2020; Song et al., 2021a; 2023; Rombach et al., 2022), and autoencoding (Kingma & Welling, 2013; He et al., 2022; Li et al., 2023) objectives. Many systems in the generative and autoencoding paradigms rely on “encoder-decoder” architectures, often styled after the original U-Net (Ronneberger et al., 2015), which contains additional long-range shortcuts between corresponding layers in mirrored symmetry about a central bottleneck. With representation learning as a goal, one typically desires that the middle bottleneck layer produce a feature embedding reflecting abstract semantic properties. The interaction of skip-connection scaffolding for gradient propagation with encoder-decoder architectures, self-supervised training objectives, and bottleneck representations has not been carefully reconsidered. This is a worrisome oversight, especially since, even in the supervised setting with standard classification architectures, prior work suggests that unweighted identity shortcuts may be a suboptimal design decision (Savarese & Figueiredo, 2017; Fischer et al., 2023).
@@
Intuitively, identity shortcut connections may not be entirely appropriate for capturing high-level, semantic representations as they directly inject low-level, high-frequency details of inputs into outputs, potentially compromising feature abstraction. We explore this issue within generative learning frameworks, including masked autoencoders (MAEs) (He et al., 2022) and diffusion models (Ho et al., 2020), leading paradigms for self-supervised image representation learning and generation. Our experiments demonstrate that identity shortcuts significantly harm semantic feature learning in comparison to an alternative we propose: gradually decay the weight of the identity shortcut over the depth of the network, thereby reducing information flow through it (Figure 2). With increasing layer depth, our approach facilitates a smooth transition from a residual to a feed-forward architecture, while maintaining sufficient connectivity to train the network effectively. Unlike prior work on learned gating (Srivastava et al., 2015a) or reweighting (Savarese & Figueiredo, 2017) mechanisms for residual connections, our method is a forced decay scheme governed by a single hyperparameter.
@@
A parallel motivation for our design stems from Huh et al. (2021), who show that features from residual blocks have higher rank than those produced by comparative feed-forward blocks. The smooth transition between residual and feed-forward behavior induced by our decay scheme regularizes deeper features toward exhibiting low-rank characteristics. Section 6 experimentally explores the correlation between our decayed identity shortcuts and low-rank feature representations. Figure 1 previews the corresponding improvements to representation learning. Our contributions are:
@@
    We introduce decayed identity shortcuts, a simple architectural mechanism which enhances semantic feature abstraction in masked autoencoders and diffusion models.
@@
    Our novel design within an MAE yields a substantial performance boost in linear probing on ImageNet-1K (Deng et al., 2009) (72.7% from a baseline 67.8%).
@@
    In diffusion models, our design improves both feature learning and generation quality.
  • $\bullet$
    Ablation studies on ImageNet-100 show that smaller models equipped with decayed identity shortcuts outperform larger ones using standard residual connections. A VIT-S/16 model (Dosovitskiy et al., 2021) with our shortcut design outperforms a baseline VIT-B/16 (78.5% vs. 76.5%).
@@ 2 Related Work
Self-supervised representation learning. Recent advancements (Achiam et al., 2023; Kirillov et al., 2023; Rombach et al., 2022; Team et al., 2023; Shi et al., 2020; Ramesh et al., 2021) in deep learning follow a common scaling law, in which a model’s performance consistently improves with its capacity and the size of the training data. This effect can be observed in large language models (LLMs), which are trained on vast amounts of internet text, enabling them to perform some tasks at human level (Laskar et al., 2023) and exhibit remarkable zero-shot capabilities (Kojima et al., 2022). These models are trained using next-token-prediction, allowing them to be trained without labeled data. In contrast, the progress of this scaling law in computer vision has largely depended on annotated data. For instance, the Segment Anything model (Kirillov et al., 2023) leverages 1 billion human-annotated masks, and state-of-the-art image generators (Ramesh et al., 2021) require training on huge datasets of text-image pairs (Schuhmann et al., 2022). However, the vast volume of unlabeled visual data and desire for continued scaling motivates a transition to self-supervised learning.
@@
At present, two families of approaches to self-supervised visual representation learning appear particularly promising: contrastive learning (Wu et al., 2018; He et al., 2020; Chen et al., 2020; Caron et al., 2021; Grill et al., 2020), which trains a discriminative model to maximize mutual information across image augmentations, and generative learning, via masked image modeling (Bao et al., 2022; He et al., 2022; Chen et al., 2024), which trains to reconstruct occluded pixels, or via diffusion denoising (Song et al., 2021b; Ho et al., 2020; Song et al., 2021a), which trains to reverse a process that mixes images with Gausssian noise. Some hybrid approaches (Zhou et al., 2022; Huang et al., 2023; Li et al., 2023) combine both families. Despite advancements, neither has demonstrated the same scalability (Singh et al., 2023) as seen in LLMs. This challenge is additional motivation for reconsidering the foundations of self-supervised network architectures.
@@
Residual and skip-connection architectures. Highway networks (Greff et al., 2017) first propose an additive skip connection structure to provide a scaffolding for gradient propagation when training very deep (e.g., 100 layer) networks. Motivated by the gating mechanisms within LSTMs (Hochreiter & Schmidhuber, 1997), this solution uses learned gating functions to weight each combination of identity and layer output branches. Residual networks (He et al., 2016) are a simplification that removes these learned coefficients. DenseNet (Huang et al., 2017) and FractalNet (Larsson et al., 2017) demonstrate that access to gradient paths of multiple lengths are the core requirement of training scaffolding, by introducing skip-connection structures with other functional forms. DenseNet utilizes feature concatenation instead of addition, while FractalNet imposes a recursive tree-like architecture combining subnetworks of multiple depths.
@@
Zhu et al. (2018) explore variants of ResNets and DenseNets with fewer points of combination between different internal paths, demonstrating that a sparser scaffolding structure may be more robust as network depth increases to thousands of layers. Savarese & Figueiredo (2017) add a scalar gating functional to the layer output in residual networks, yielding a hybrid design between residual and highway networks; learning this scalar gating provides a consistent benefit to classification accuracy. Fischer et al. (2023) develop a weighting scheme for residual connections based upon a sensitivity analysis of signal propagation within a ResNet. To date, none of these potential improvements have seen broad adoption.
@@
Low rank bias in neural networks. Over-parameterized neural networks exhibit surprising generalization capabilities, a finding seemingly in contradiction with classical machine learning theory (Neyshabur et al., 2019). This phenomenon implies the existence of some form of implicit regularization that prevents the model from overfitting. From the perspective of neural network parameterizations, Arora et al. (2019) suggest that linear models with more layers tend to converge to minimal norm solutions. In the context of CNNs, Huh et al. (2021) demonstrates that stacking more feed-forward layers compels the model to seek solutions of a lower rank, and Jing et al. (2020) reinforce this finding by adding more layers to an autoencoder’s bottleneck, thereby creating a representation bottleneck. In vision transformers, Geshkovski et al. (2024) examine the connection between attention blocks and mean-shift clustering (Cheng, 1995), showing that repeated attention operations result in low-rank outputs. Moreover, Dong et al. (2021) reveal that eliminating the shortcut connection from residual attention blocks causes features to degenerate to rank 1 structures doubly exponentially. From a different perspective, recent work (Radhakrishnan et al., 2022; Beaglehole et al., 2023; Radhakrishnan et al., 2024) shows training algorithms implicitly induce low-rank behavior in neural networks. Radhakrishnan et al. (2024) study the dimensionality reduction behavior of a recursive feature machine (Radhakrishnan et al., 2022) and effectively verify performance on low-rank matrix recovery.
@@ 3 Method
Prior works show that deeper feed-forward architectures have an inductive bias towards producing low-rank feature maps, while ResNets do not display the same behavior (Huh et al., 2021). However, despite this bias, deeper feed-forward architectures are typically less effective and generalize worse than ResNets (He et al., 2016). We aim to combine the properties of both feed-forward networks and ResNets, using the low-rank prior to enhance the abstraction capability of the network while maintaining the core benefits of the residual block, including stable training and the capacity to construct deeper models.
@@ 3.1 Decayed Identity Shortcuts
where ${\bm{x}}_{l}\in\mathbb{R}^{d}$ represents the output from the preceding layer, and $f_{{\bm{\theta}}_{l}}$ denotes the transformation applied at the current layer. Although it is widely known that pure feed-forward architectures are susceptible to vanishing gradients when building deeper models, Huh et al. (2021) demonstrates that feed-forward modules offer implicit structural regularization, enabling deep models to generate abstract representations at bottlenecks.
Residual connections. To address the optimization problem of vanishing gradients in deeper neural networks, ResNets (He et al., 2016) construct each layer as a residual function, resulting in a modification to Eq. 1:
@@
This design builds shortcuts from input to output, allowing gradient magnitude to be preserved regardless of the depth of the model. However, a consequence of this design is that the output stays close to the input in practice (Greff et al., 2017), defeating the need to construct complex transformations over depth. The same phenomenon is also observed in highway networks (Srivastava et al., 2015a), which adopt learnable gates $H_{\phi}({\bm{x}})\in[0,1]^{d}$ in both the residual and skip branches: ${\bm{x}}_{l+1}=H_{\phi}({\bm{x}}_{l})\cdot{\bm{x}}_{l}+(1-H_{\phi}({\bm{x}}_{l}))\cdot f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})$. Although this flexible design allows the model to build the abstraction level over depths, similar to feedforward networks, Srivastava et al. (2015b) finds $H_{\phi}\approx 1$ for most units, suggesting the model prefers copying the input.
Decayed identity shortcuts for unsupervised representation learning. Setting aside the optimization benefits brought by residual connections, we rethink the role of the residual connections from the viewpoint of representation learning. Abstraction can be viewed as invariance to local changes of input and is crucial to the disentanglement of the feature space (Bengio et al., 2013). Prior work suggests that a shortcut path of residual connections tends to preserve high-frequency fine-grained input information (Greff et al., 2017), resulting in decreased feature abstraction. We hypothesize that this lack of abstraction harms the capability of the model to learn meaningful low-level features and that ensuring an abstract structure in the deeper layers of the neural network will help improve representation learning, especially for unsupervised tasks that often use indirect proxy objectives, such as pixel-wise reconstruction loss. Motivated by this hypothesis, we propose to downweight the contribution from the shortcut path: